# Latent Schrödinger Bridge Diffusion Model for Generative Learning

### v1

@@ 4.2.1 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widehat{\pi}_{T}},\pi_{T})]$
denote the statistical error and approximation error, respectively. Leveraging the tools of empirical process theory [VW23, VdV00] and deep approximation theory [LSYZ21, PV18, Yar17, CLZZ20, JSLH23], we can bound these errors, as expounded in the following two lemmas.
@@
Suppose that Assumption 4.3 holds. Given an approximation error $\epsilon>0$ and $T\in(0,1)$, we can choose a neural network $\mathbf{s}$ with the following structure:
$L=\mathcal{O}\left(\log{\frac{1}{\epsilon}}+d^{*}\right),M=\mathcal{O}\left(\frac{{d^{*}}^{\frac{5}{2}}\left(\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{d^{*}+3}{2}}{\xi}^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{3}}\right),$
$J=\mathcal{O}\left(\frac{{d^{*}}^{\frac{5}{2}}\left(\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{d^{*}+3}{2}}\xi^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{3}}\left(\log{\frac{1}{\epsilon}}+d^{*}\right)\right),K=\mathcal{O}\left(\frac{\sqrt{d^{*}\log{\frac{d^{*}}{\epsilon(1-T)}}}}{1-T}\right),$
$\kappa=\mathcal{O}\left(\xi\sqrt{\log{\frac{d^{*}}{\epsilon(1-T)}}}\vee\frac{\left(d^{*}\log\frac{d^{*}}{\epsilon(1-T)}\right)^{\frac{3}{2}}}{(1-T)^{3}}\right),$
such that for any $t\in[0,T]$, we have
$\mathbb{E}_{\mathbf{x}_{t}}\|\mathbf{s}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\leq(1+d^{*})\epsilon^{2},$
where $\xi$ is the Lipschitz constant for $\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x})$ w.r.t. $\mathbf{x}$.
@@
Given an approximation error $\epsilon>0$ and $T\in(0,1)$, the score estimator $\widehat{\mathbf{s}}$ defined in (6), with the neural network structure introduced in Lemma 4.4, satisfies
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)=\widetilde{\mathcal{O}}\left(\frac{1}{n}\cdot\frac{\xi^{d^{*}}\epsilon^{-(d^{*}+1)}}{(1-T)^{5}}\right),$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}(\widehat{\mathbf{s}})\right)=\widetilde{\mathcal{O}}\left(\frac{1}{\sqrt{m}}\cdot\frac{\xi^{\frac{d}{2}}\epsilon^{-\frac{d^{*}+1}{2}}}{(1-T)^{\frac{7}{2}}}\right).$
By combining the approximation and statistical errors presented in Lemmas 4.4-4.5, we can now derive the upper bound for the score estimation, as detailed in the following theorem.
@@
Suppose Assumption 4.3 holds and let $T\in(0,1)$. By choosing $\epsilon=n^{-\frac{1}{d^{*}+3}}$ in Lemmas 4.4-4.5, the score estimator $\hat{s}$ defined in (6) satisfies
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(t,\mathbf{x}_{t})-\nabla\log{q}(\sqrt{1-t}\sigma,\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)$
$=\widetilde{\mathcal{O}}\left(\frac{\xi^{d^{*}}}{(1-T)^{5}}\left(n^{-\frac{2}{d^{*}+3}}+n^{\frac{d^{*}+1}{2(d^{*}+3)}}m^{-\frac{1}{2}}\right)\right).$
@@
In the theoretical analysis of existing diffusion models, a substantial number of inquiries prominently espouse the $L_{2}$-accuracy of the score estimator. This is evident in a spectrum of assumptions articulated by the literature, including [CLL23, Assumption 1], [CDS23, Assumption H1], [LLT22, Assumption 2], [LLT23, Assumption 1], [BDBDD23, Assumption 1], [LWCC23, Assumption 1], and [GNZ23, Assumption 3]. Despite this pervasive reliance on such assumptions, there exists a dearth of research endeavors that have undertaken a rigorous theoretical consideration of score estimation. Noteworthy among the limited investigations in this domain is [WJX+21, Theorem 5], which scrutinizes the convergence consistency of score estimation. Furthermore, both [OAS23, Theorem 3.1] and [CHZW23, Theorem 2] contribute to the field by providing the convergence rate of score estimation. It is imperative to underscore that [OAS23, CHZW23] operated within a continuous-time paradigm, neglecting considerations of time discretization and presuming the feasibility of resolving integrals over the time interval $[0,T]$. In contradistinction, our results meticulously incorporate considerations of time discretization, offering a convergence rate within the discretized time.
With Theorem 4.6, we can bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\pi_{T})]$ as the following theorem.
@@
Suppose that assumptions of Theorem 4.6 hold, then we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\pi_{T})]=\widetilde{\mathcal{O}}\left(\frac{e^{\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\xi^{\frac{d^{*}}{2}}}{(1-T)^{\frac{5}{2}}}\left(n^{-\frac{1}{d^{*}+3}}+n^{\frac{d^{*}+1}{4(d^{*}+3)}}m^{-\frac{1}{4}}\right)\right),$
where $\gamma_{1}=10d^{*}\xi$.
4.2.2 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\widetilde{\pi}_{T})]$
Recall that the distribution of $\widetilde{\mathbf{x}}_{t}$ is denoted by $\widetilde{\pi}_{t}$. The error bound between $\widehat{\pi}_{T}$ and $\widetilde{\pi}_{T}$ stems from the EM discretization. By employing numerical techniques of SDE, this error can be characterized through the elucidation provided in the following theorem.
Theorem 4.8 (Error Bound for EM Discretization).
Suppose that Assumption 4.3 holds. Given $T\in(0,1)$ and an approximation error $\epsilon>0$, let $\{t_{k}\}_{k=0}^{N-1}$ be the partition for $[0,T]$, then we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}(\widehat{\pi}_{T},\widetilde{\pi}_{T})]=\widetilde{\mathcal{O}}\left(\left(\gamma_{1}K+\gamma_{2}\right)e^{\sigma^{2}\sqrt{d^{*}}\gamma_{1}}\sqrt{\sum_{k=0}^{N-1}(t_{k+1}-t_{k})^{3}}\right),$
where $\gamma_{1}=10d^{*}\xi$ and $\gamma_{2}=\mathcal{O}\left(\frac{10\left(d^{*}\log{\frac{d^{*}}{\epsilon(1-T)}}\right)^{\frac{3}{2}}}{(1-T)^{3}}\right)$.
4.2.3 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$
In this section, we present the upper bound for $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$. This error originates from the process of early stopping. From the definition of $\pi_{T}$, it is observed that $\pi_{T}$ converges to $p_{data}$ as $T$ approaches 1. Consequently, the characterization of $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T},{\widehat{p}}^{*}_{data})]$ can be expressed in terms of $T$, elucidated in the following lemma.

### v2

@@ 4.2 Error Bounds in Latent Diffusion Modeling
denote the statistical error and approximation error, respectively.
Using the tools of empirical process theory and deep approximation theory, we can bound these errors. We first introduce three necessary assumptions.
Assumption 4.3 (Bounded Density).
The density $\widehat{p}_{data}^{*}$ in the latent space is supported on $[-1,1]^{d^{*}}$ and is bounded above and below by $C_{u}$ and $C_{l}$, respectively, on $[-1,1]^{d^{*}}$.
Assumption 4.4 (Hölder Continuous).
The density $\widehat{p}_{data}^{*}$ in the latent space belongs to $\mathcal{H}^{\beta}([-1,1]^{d^{*}},R)$ for a Hölder index $\beta>0$ and constant $R>0$.
Assumption 4.5 (Boundary Smoothness).
There exists $10<a_{0}<$ 1 such that $\widehat{p}_{data}^{*}\in C^{\infty}([-1,1]^{d^{*}}\backslash[-1+a_{0},1-a_{0}]^{d^{*}})$.
Assumption 4.3 introduces the compact support for the target distribution in the latent space, which is similarly adopted in [LLT23, LWCC23, OAS23]. Assumption 4.4 enforces the Hölder continuity of the density, a fundamental condition in nonparametric estimation. Assumption 4.5 imposes a regularity condition on the density $\widehat{p}_{data}^{*}$, a technical assumption that is also introduced in [OAS23]. In the existing literature, certain regularity conditions on the target distribution have also been introduced. For instance, it is required that the gradient of the score function be bounded both from above and below. Such conditions are outlined in [LLT22, Assumption 1], [CLL23, Assumptions 3-4], [LLT23, Assumptions 3-4], [GNZ23, Assumption 1], [OAS23, Assumptions 2.4, 2.6], and [CHZW23, Assumption 3]. Moreover, [CHZW23, Assumptions 1, 2] introduced analogous conditions. Their Assumption 1 delineates the representation of data points $\mathbf{x}\sim p_{data}$ as a product of an unknown matrix $A$ with orthonormal columns and a latent variable $\mathbf{z}$ following a distribution $P_{z}$ with a density function $p_{z}$. Meanwhile, their Assumption 2 supposes the density function $p_{z}>0$ is twice continuously differentiable. Moreover, there exist positive constants $B,C_{1},C_{2}$ such that when $\|\mathbf{z}\|_{2}\geq B$, the density function $p_{z}(\mathbf{z})\leq$ $(2\pi)^{-d/2}C_{1}\exp\left(-C_{2}\|\mathbf{z}\|_{2}^{2}/2\right)$.
Under Assumptions 4.3-4.5, we provide the following lemma, which bounds the approximation error.
@@
Suppose that Assumptions 4.3-4.5 hold. Let $\mathcal{R}\gg 1$, $C_{T}>0$, and $T=1-\mathcal{R}^{-C_{T}}$. We can then choose a ReLU neural network $\mathbf{s}\in\mathrm{NN}(L,M,J,\kappa)$ that satisfies
$\mathbf{s}(t,\mathbf{x})=\mathbf{s}(t,\mathbf{s}_{\mathrm{clip}}(\mathbf{x},-C_{0},C_{0}))\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \|\mathbf{x}\|_{\infty}>C_{0},\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ C_{0}=\mathcal{O}(\sqrt{\log\mathcal{R}}),$
and has the following structure:
$L=\mathcal{O}(\log^{4}\mathcal{R}),M=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{7}\mathcal{R}),J=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{9}\mathcal{R}),\kappa=\exp\left(\mathcal{O}(\log^{4}\mathcal{R})\right).$
Then, for any $t\in[\mathcal{R}^{-C_{T}},1]$, we have
$\int_{\mathbf{x}\sim q_{t}(\mathbf{x})}\|\mathbf{s}(t,\mathbf{x})-\nabla\log q_{t}(\mathbf{x})\|^{2}\mathrm{d}\mathbf{x}\lesssim\frac{\mathcal{R}^{-2\beta}\log\mathcal{R}}{\sigma_{t}^{2}}.$
Moreover, we can take $\mathbf{s}$ satisfying $\|\mathbf{s}(t,\cdot)\|_{\infty}\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}}$.
Now, we restrict the ReLU neural network class $\mathrm{NN}(L,M,J,\kappa)$ to
$\mathcal{C}:=\Big{\{}\mathbf{s}\in$ $\mathrm{NN}(L,M,J,\kappa)\Big{|}\|\mathbf{s}(t,\cdot)\|_{\infty}\lesssim\frac{\sqrt{\log\mathcal{R}}}{\sigma_{t}},$
$\mathbf{s}(t,\mathbf{x})=\mathbf{s}(t,\mathbf{s}_{\mathrm{clip}}(\mathbf{x},-C_{0},C_{0})\leavevmode\nobreak\ \text{for}\leavevmode\nobreak\ \|\mathbf{x}\|_{\infty}>C_{0},\leavevmode\nobreak\ \text{where}\leavevmode\nobreak\ C_{0}=\mathcal{O}(\sqrt{\log\mathcal{R}}),$
$L=\mathcal{O}(\log^{4}\mathcal{R}),M=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{7}\mathcal{R}),J=\mathcal{O}(\mathcal{R}^{d^{*}}\log^{9}\mathcal{R}),\kappa=\exp\left(\mathcal{O}(\log^{4}\mathcal{R})\right)\Big{\}}.$
We note that the method for deriving the upper bound on the approximation error in Lemma 4.3 is inspired by [OAS23, FYWC24]. To elaborate, we express the score function $\nabla\log q_{t}(\mathbf{x})$ as
$\nabla\log q_{t}(\mathbf{x})=\frac{\nabla q_{t}(\mathbf{x})}{q_{t}(\mathbf{x})}.$
To obtain the approximation, we treat the numerator and denominator separately. Since the approximation techniques for both terms are analogous, we focus on approximating $q_{t}(\mathbf{x})$ as a representative example. Further details on the full approximation procedure are provided in Section B of the appendix. The first step involves truncating the integral domain in $q_{t}(\mathbf{x})$. Recall that
$q_{t}(\mathbf{x})=\int_{\mathbb{R}^{d^{*}}}\widehat{p}_{data}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\underbrace{\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)}_{\text{Transition Kernel}}\mathrm{d}\mathbf{y}.$
There exists a constant $C>0$ such that for any $\mathbf{x}\in\mathbb{R}^{d^{*}}$,
$\left|q_{t}(\mathbf{x})-\int_{A_{\mathbf{x}}}\widehat{p}_{data}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)\mathrm{d}\mathbf{y}\right|\lesssim\epsilon,$
where $A_{\mathbf{x}}=\prod_{i=1}^{d^{*}}a_{i,\mathbf{x}}$ with $a_{i,\mathbf{x}}=\left[x_{i}-C\sigma_{t}\sqrt{\log\epsilon^{-1}},x_{i}+C\sigma_{t}\sqrt{\log\epsilon^{-1}}\right]$. Next, we approximate the integral over $A_{\mathbf{x}}$. To do so, we introduce a Taylor polynomial $f_{Taylor}^{*}(\mathbf{y})$ to approximate $\widehat{p}_{data}^{*}(\mathbf{y})$. As a result, $q_{t}(\mathbf{x})$ is approximated as
$\int_{A_{\mathbf{x}}}f_{Taylor}^{*}(\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\cdot\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)\mathrm{d}\mathbf{y}.$
In the third step, we approximate the exponential transition kernel $\frac{1}{\sigma_{t}^{d^{*}}(2\pi)^{d^{*}/2}}\exp\left(-\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\sigma_{t}^{2}}\right)$ using a Taylor polynomial denoted as $f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})$. This yields the approximation
$\int_{A_{\mathbf{x}}}f_{Taylor}^{*}(\mathbf{y})f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})\mathbf{I}_{\{\|\mathbf{y}\|_{\infty}\leq 1\}}\mathrm{d}\mathbf{y}.$
Since the product $f_{Taylor}^{*}(\mathbf{y})f_{Taylor}^{kernel}(t,\mathbf{x},\mathbf{y})$ is a polynomial, its integration can be evaluated explicitly. In the fourth step, we utilize a deep neural network to approximate the resulting integral. Integrating these analyses, we can ultimately construct a ReLU neural network that approximates the score function.
@@
Suppose that Assumptions 4.3-4.5 hold. Let $\mathcal{R}\gg 1$, $C_{T}>0$, and $T=1-\mathcal{R}^{-C_{T}}$. The score estimator $\widehat{\mathbf{s}}$ defined in (6), with the neural network structures belonging to class $\mathcal{C}$, satisfies
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\mathcal{L}(\widehat{\mathbf{s}})-2\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})+\mathcal{L}(\mathbf{s}^{*})\right)\lesssim\frac{\mathcal{R}^{d^{*}}\log^{15}\mathcal{R}\left(\log^{4}\mathcal{R}+\log n\right)}{n}$
@@
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\overline{\mathcal{L}}_{\widehat{\mathcal{X}}}(\widehat{\mathbf{s}})-\widehat{\mathcal{L}}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}(\widehat{\mathbf{s}})\right)\lesssim\mathcal{R}^{C_{T}}(\log m+\log\mathcal{R})\cdot\frac{\mathcal{R}^{\frac{d^{*}}{2}}\log^{\frac{13}{2}}\mathcal{R}(\log^{2}\mathcal{R}+\sqrt{\log m})}{\sqrt{m}}.$
By combining the approximation and statistical errors presented in Lemmas 4.3-4.4, we can now derive the upper bound for the score estimation, as detailed in the following theorem.
@@
Suppose that Assumptions 4.3-4.5 hold. By choosing $\mathcal{R}=\lfloor n^{\frac{1}{d^{*}+2\beta}}\rfloor+1\lesssim n^{\frac{1}{d^{*}+2\beta}}$, $m=n^{\frac{d^{*}+8\beta}{d^{*}+2\beta}}$, and $C_{T}=2\beta$ in Lemmas 4.3-4.4, the score estimator $\widehat{\mathbf{s}}$ defined in (6), with the neural network structure belonging to class $\mathcal{C}$, satisfies
$\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left(\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(1-t,\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right)\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n.$
@@
In the theoretical analysis of existing diffusion models, a substantial number of inquiries prominently espouse the $L_{2}$-accuracy of the score estimator. This is evident in a spectrum of assumptions articulated by the literature, including [CLL23, Assumption 1], [CDS23, Assumption H1], [LLT22, Assumption 2], [LLT23, Assumption 1], [BDBDD23, Assumption 1], [LWCC23, Assumption 1], and [GNZ23, Assumption 3]. Despite this pervasive reliance on such assumptions, there exists a dearth of research endeavors that have undertaken a rigorous theoretical consideration of score estimation. Noteworthy among the limited investigations in this domain is [WJX+21, Theorem 5], which scrutinizes the convergence consistency of score estimation. Furthermore, both [OAS23, Theorem 3.1] and [CHZW23, Theorem 2] contribute to the field by providing the convergence rate of score estimation. It is imperative to underscore that [OAS23, CHZW23] operated within a continuous-time paradigm, neglecting considerations of time discretization and presuming the feasibility of resolving integrals over the time interval $[0,T]$. In contrast, our results meticulously incorporate considerations of time discretization, offering a convergence rate within the discretized time. Importantly, by ignoring the logarithmic term, this bound achieves the minimax optimal rate in nonparametric statistics [Sto80, Tsy09].
Now, we consider bounding $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\widehat{p}^{*}_{data})]$. We systematically decompose this error term into two distinct components, as outlined by the following inequality:
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\widehat{p}^{*}_{data})]\leq\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\pi_{T}^{L})]+\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},\widehat{p}^{*}_{data})].$ (9)
Subsequently, we proceed to establish individual bounds for the two terms on the right-hand side of (9) in Subsections 4.2.1-4.2.2. The integration of these two analyses culminates in the oracle inequality for $\mathbb{E}_{\mathcal{X},\mathcal{Y}}[W_{2}({\widetilde{\pi}_{T}^{L}},\widehat{p}_{data}^{*})]$, which is expounded upon in Subsection 4.2.3.
4.2.1 Bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\pi_{T}^{L})]$
In this subsection, our aim is to establish an upper bound for $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}({\widetilde{\pi}_{T}^{L}},\pi_{T}^{L})]$. With Theorem 4.5, for any $\epsilon>0$, there exists $\Delta_{\epsilon}>0$ such that if $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})\leq\Delta_{\epsilon}$, then
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbf{x}_{t_{i}}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})\|^{2}\mathrm{d}t\right]$
$=\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbf{x}_{t_{i}}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})\|^{2}\mathrm{d}t\right]$
$\leq T\cdot\mathbb{E}_{\widehat{\mathcal{X}},\mathcal{T},\mathcal{Z}}\left[\frac{1}{T}\int_{0}^{T}\mathbb{E}_{\mathbf{x}_{t}}\|\widehat{\mathbf{s}}(1-t,\mathbf{x}_{t})-\nabla\log q_{1-t}(\mathbf{x}_{t})\|^{2}\mathrm{d}t\right]+\epsilon\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n+\epsilon.$
Specially, taking $\epsilon=n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n$, there exists $\Delta_{\epsilon}=\Delta_{n}$ such that if $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})\leq\Delta_{n}$, then we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}\left[\sum_{i=0}^{K-1}\int_{t_{i}}^{t_{i+1}}\mathbb{E}_{\mathbf{x}_{t_{i}}}\|\widehat{\mathbf{s}}(1-t_{i},\mathbf{x}_{t_{i}})-\nabla\log q_{1-t_{i}}(\mathbf{x}_{t_{i}})\|^{2}\mathrm{d}t\right]\lesssim n^{-\frac{2\beta}{d^{*}+2\beta}}\log^{19}n.$ (10)
Using (10), we can bound $\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]$ as the following theorem.
@@
Suppose that assumptions of Theorem 4.5 hold. By choosing $\max_{0\leq i\leq K-1}(t_{i+1}-t_{i})=\mathcal{O}\left(\min\{\Delta_{n},n^{-\frac{6\beta}{d^{*}+2\beta}}\}\right)$, then we have
$\mathbb{E}_{\mathcal{X},\mathcal{Y},\mathcal{T},\mathcal{Z}}[W_{2}(\widetilde{\pi}_{T}^{L},\pi_{T}^{L})]\lesssim n^{-\frac{\beta}{d^{*}+2\beta}}\log^{\frac{19}{2}}n.$
4.2.2 Bound $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},{\widehat{p}}^{*}_{data})]$
In this subsection, we present the upper bound for $\mathbb{E}_{\mathcal{Y}}[W_{2}(\pi_{T}^{L},{\widehat{p}}^{*}_{data})]$. This error originates from the process of early stopping. We have the following lemma.